OpenStudy (anonymous):

Which is the equation to the line passing through points (2, –7) and (5, –1)?

OpenStudy (netlopes1):

If you know this equation (to distance between two points), so it's easier:\[d_{A},_{B}=\sqrt{(x_a-x_b)^2+(y_a-y_b)^2}\] with \[A(x_a,y_a)\] and \[B(x_b,y_b)\]

OpenStudy (netlopes1):

Make this: Call A(2, -7) and B(5, -1) and put this values in the equation....

OpenStudy (netlopes1):

Sorry....... Now i see...... you want only equation....

OpenStudy (netlopes1):

What i did is correct, but you don't want this.... sorry again.... Give me try again ok?

OpenStudy (anonymous):

y = mx + b where m = slope, b = constant use y2-y1/x2-x1 to find m. sub that in with either one of those points to find b. done.

OpenStudy (netlopes1):

Still call A(2, -7) and B(5, -1) and use this.... yes, "petewe", you are correct....

OpenStudy (netlopes1):

I'll use, A(2, -7) and B(5, -1) to find "m" (slope). So \[m=[-1-(-7)]/[5-2]\rightarrow m=6/3\rightarrow m=2\]

OpenStudy (netlopes1):

Now to find "b", do y=2x+b with only one point, "A" or "B". I use A(2, -7) and \[-7=2(2)+b \rightarrow -7=4+b \rightarrow b=-11\]

OpenStudy (netlopes1):

Finally, the equation is: y = 2x - 11. Ok?

OpenStudy (anonymous):

thank you so much